# Must Read: A Systematic Survey of Computational Persuasion

### v1

@@ 3.3. LLM Persuasiveness
Unlike the evaluation of argument persuasiveness, where arguments are predefined by humans, LLM persuasiveness evaluation focuses on model-generated content in both single-turn and multi-turn dialogues. In this survey, we refer to this setting as AI as Persuader, where the language model acts as the source of persuasion, aiming to influence a human’s beliefs, attitudes, or behaviors through generated arguments or dialogue. Leading companies such as OpenAI (OpenAI, 2024), Anthropic (Durmus et al., 2024), and DeepMind (Phuong et al., 2024) have investigated the persuasiveness of their models to better understand the risks associated with deploying these highly conversational systems. However, evaluating the persuasive abilities of LLMs is not straightforward, as research explores a variety of evaluation methodologies and experimental setups.
@@ 3.3.1. Human Evaluation
Understanding the dynamics of AI to Human persuasion is crucial, as most proprietary models are designed for individual users. Given this, using human subjects to assess the persuasiveness of LLMs is a natural approach.
@@
Durmus et al. (2024) assessed model persuasiveness by measuring human agreement with a claim before and after exposure to persuasive arguments generated by Claude models. This single-turn setup reveals that larger models are generally more persuasive, with deceptive techniques emerging as particularly effective. In contrast, Phuong et al. (2024) introduced a suite of multi-turn evaluation benchmarks, including Money Talks, Charm Offensive, Hidden Agenda, and Web of Lies. These involve interactive persuasion tasks such as persuading users to donate, impersonating a friend, manipulating users to take suspicious actions on the computer, and promoting false beliefs. Unlike single-turn approaches, these multi-turn evaluations are expected to capture more dynamic and complex forms of persuasion.
@@
Similarly, Salvi et al. (2024) evaluated model persuasiveness through a debate game involving both human-human and human-LLM interactions. They found that participants were more likely to change their stance after engaging with GPT-4, regardless of whether the model had access to personal information. This suggests that GPT-4 was more persuasive than human debaters across conditions. OpenAI (2024) introduced two human evaluation benchmarks for assessing LLM persuasiveness. In the first setup, they used successful persuasion samples from the r/ChangeMyView subreddit as prompts, then asked LLMs to generate persuasive arguments in response. Human annotators scored both the original human arguments and the LLM-generated ones. The primary evaluation metric, referred to as the AI persuasiveness percentile relative to humans, measures the probability that a randomly chosen LLM-generated argument is rated as more persuasive than its human counterpart. The second benchmark, called Persuasion Parallel Generation, presents annotators with two arguments generated by different models for the same prompt. Annotators select the more persuasive argument, and the resulting win rates are used to compare model performance. Results from both benchmarks indicate that LLMs do not significantly outperform humans, and that newer models such as o1 do not show substantial gains in persuasiveness over earlier generations.
@@
Interestingly, similar lines of research have emerged within the social sciences. Huang and Wang (2023) conducted a meta-analysis examining whether artificial intelligence is more persuasive than humans. Using statistical methods, they compared persuasion outcomes—such as changes in perception, attitudes, intentions, and behaviors—between human-AI and human-human interactions. Their findings revealed a small overall Cohen’s d (Cohen, 1988), suggesting that LLMs are about as persuasive as humans. Focusing more specifically on political communication, Goldstein et al. (2024) found that AI-generated propaganda is already as persuasive as real-world examples. Similarly, Bai et al. (2025) showed that AI performs on par with humans in crafting messages aimed at shaping public attitudes toward policy issues.
@@ 3.3.2. Automatic Evaluation
Singh et al. (2024) proposed PersuasionBench and PersuasionArena to evaluate LLMs’ persuasive abilities using tweet pairs from enterprise accounts that differ in engagement. They introduced simulative tasks (e.g., predicting or comparing tweet persuasiveness) and generative transsuasion tasks that modify content to improve engagement. Their evaluations spanned standard NLP metrics, human judgments, and an Oracle LLaMA-13B model. While automated metrics and human or Oracle evaluations do not always align, they generally support the conclusion that the models proposed by Singh et al. (2024) achieve the strongest performance. Another work in this direction, Pauli et al. (2025), trained a regression model to score the relative persuasiveness of LLM-generated rewrites, using human-annotated text pairs collected across diverse domains.
@@
Building on the conversational persuasion evaluation paradigm, OpenAI (2024) proposed domain-specific scenarios to assess persuasive abilities. In MakeMeSay, the persuader is given a secret codeword and must steer the persuadee into saying it; in MakeMePay, the persuader adopts the role of a con artist attempting to elicit a payment. Similarly, Zhu et al. (2025) introduced a multi-agent bargaining task, where groups of LLM sellers attempt to persuade LLM buyers to purchase a product. Idziejczak et al. (2025) explored persuasion in a social deception game inspired by Among Us, where LLMs played as either crewmates or impostors. Persuasive behavior emerged through dialogue, which was later annotated by another LLM using predefined persuasion strategies to assess how models influenced one another within the game.
Although automatic approaches to LLM persuasiveness evaluation help facilitate scalability, making it possible to evaluate persuasion across the rapid iteration of new LLMs and avoid overhead related to difficulty in designing human-subject research, they still suffer from contradictory evaluation results across different test cases. For instance, in Bozdag et al. (2025), GPT models achieved a margin of persuasiveness over Claude models, while in the test cases used by Idziejczak et al. (2025), Claude models were performing better than GPT models.
While early findings suggest that LLMs can be as persuasive as humans—and in some cases more so, particularly as they improve in reasoning, personalization, and strategic communication, the underlying mechanisms driving this persuasiveness remain poorly understood. Current research has revealed that different evaluation setups may yield inconsistent or even contradictory results, with some models outperforming others in one framework while falling short in another. This highlights the complex, multidimensional nature of persuasive ability and raises important concerns about the increasing influence of ”smarter” models on beliefs and behaviors. As a result, developing a unified and trustworthy evaluation framework that integrates diverse test cases and assesses persuasive skills across multiple dimensions is a pressing research challenge. Such a framework would not only support more self-consistent and interpretable evaluations but also help explain why certain models excel in specific contexts while lacking in others, ultimately informing safer and more responsible deployment of persuasive AI systems.
@@ 4. Generating Persuasion
Systems that use an AI as Persuader to generate persuasive content are becoming increasingly more powerful and more widespread. Persuasion can play a pivotal role in various domains, including advertisements, healthcare promotion, recommendation systems, political campaigns, and more. As organizations and individuals seek to increase their influence, the need for generating persuasive content has grown significantly, especially with the advent of LLMs. This section explores prior work on enhancing persuasive capabilities, examines key influencing factors, and highlights applications of persuasion in settings like negotiation and debate.
@@ 4.1. Methods
Section 2.2.1 introduced taxonomies of persuasive strategies that LLMs can adopt to enhance the effectiveness of their outputs. Building on this foundation, recent work has focused on developing methods to improve LLMs’ ability to generate persuasive content. In this subsection, we review key approaches proposed to enhance the persuasive capabilities of LLMs, including prompting techniques, incorporation of external knowledge, fine-tuning, and reinforcement learning.
@@
(Tiwari et al., 2022b) Personalization PPMD (Tiwari et al., 2022b) Selects a persuasive strategy based on the user profile
@@
(Tiwari et al., 2022a) Personalization DevPVA (Tiwari et al., 2022a) Rewards persuasion based on persona of the user
@@ 4.1.1. Prompting
Another approach is to prompt LLMs with specific persuasive strategies. Durmus et al. (2024) examined how different types of prompts affect the persuasiveness of model outputs by experimenting with four prompting strategies: (1) prompting the model to construct compelling arguments, (2) adopting the role of an expert and using rhetorical techniques such as pathos, logos, and ethos, (3) encouraging logical reasoning, and (4) permitting deception, including fabricating evidence, to boost persuasive impact. Their findings showed that while effectiveness varies across model architectures, prompts that encourage logical reasoning or allow for deception tend to produce more persuasive arguments overall. Along similar lines, Zeng et al. (2024) designed a comprehensive taxonomy of persuasion strategies, then used these strategies to generate persuasive adversarial prompts (PAPs). This also proved to be remarkably effective in jailbreaking LLMs, outperforming alternative methods and achieving attack success rates of over 92% across Llama-2-7B-Chat, GPT-3.5, and GPT-4. Using persuasive strategies has also demonstrated effectiveness in misinformation propagation, where persuasive strategies effectively manipulate LLMs into endorsing false information (Xu et al., 2024a).
@@ 4.1.2. Incorporating External Information
While prompting offers a good baseline for persuasive generation, incorporating external information can substantially enhance persuasiveness by introducing additional dimensions such as personalization and factual grounding. Models can produce more tailored and credible persuasive content by leveraging contextual knowledge, including user profiles, background information, or external evidence.
Effective persuasion involves not only implementing persuasive strategies but also selecting techniques that align with the characteristics of the recipient. Lukin et al. (2017) and Wang et al. (2019) showed that accounting for personality traits is critical, as individuals with different personality types tend to respond differently to specific persuasive strategies. Building on this insight, Jin et al. (2024) proposed a method where LLMs begin by identifying and summarizing the user’s intent before reasoning about the most suitable strategy to apply. This process enables a more adaptive and responsive persuasive system.
Incorporating personal information about the user extends beyond just strategy selection. Personalization based on users’ psychological traits, such as personality, political ideology, moral foundations, and behavioral patterns, has been shown to significantly enhance persuasive impact (Matz et al., 2024; Kaptein et al., 2015). This approach is effective across various domains, including targeted advertisements and political messaging around climate action (Matz et al., 2024; Simchon et al., 2024; Meguellati et al., 2024). Salvi et al. (2024) showed that providing LLMs with personal information during human interaction increases their persuasiveness. Research by Ruiz-Dolz et al. (2024) demonstrated that user-specific persuasive policies significantly improve the persuasiveness of argument-based systems. Zhang and Zhou (2025) emphasized the importance of considering the user’s mental state when constructing persuasive arguments. Tiwari et al. (Tiwari et al., 2022b, 2023) showed that combining dialogue context with user profile information provides further persuasive benefits. In parallel, other research highlighted the value of counterarguments that directly engage with the user’s concerns or beliefs (Cima et al., 2024; Hunter et al., 2019).
@@ 4.1.3. Finetuning
Finetuning LLMs on persuasive datasets offers a powerful way to enhance their persuasive capabilities beyond what prompting alone can achieve. This approach allows models to internalize persuasive patterns and strategies at a more fundamental level.
In the negotiation domain, Lewis et al. (2017) trained a model that is based on four recurrent neural networks on negotiation datasets, and demonstrated significant improvements in negotiation ability. Extending finetuning to emotional support contexts, Liu et al. (2021) introduced the Emotion Support Conversation dataset (ESConv) and showed that finetuning BlenderBot (Roller et al., 2021) and DialoGPT (Zhang et al., 2020) on this data enhances models’ ability to provide empathetic responses and sustain engagement.
Chen et al. (2022) finetuned BART-large (Lewis et al., 2020) on the PersuasionForGood dataset (Wang et al., 2019), achieving significant gains in persuasive quality. More recently, Jin et al. (2024) developed PersuGPT by training LLaMa-2-13B-Chat on the DailyPersuasion dataset, which includes rich reasoning traces. Combining conventional finetuning with Direct Preference Optimization (DPO) (Rafailov et al., 2023), their model learns to infer user intent, select appropriate persuasive strategies, and anticipate future user responses and rewards. Singh et al. (2024) proposed an instruction-based finetuning approach to enhance LLM persuasiveness. They finetuned Vicuna-1.5 13B to rewrite inputs to be more persuasive and showed that the resulting model outperforms the base version and can rival much larger LLMs in persuasiveness.
Finetuning presents strong potential for developing more persuasive language models, yet remains relatively underexplored. With more carefully curated training data and domain-specific finetuning, future work can yield models that are better aligned with persuasive goals across a range of applications, including healthcare, education, and advertising.
@@ 4.1.4. Reinforcement Learning
Reinforcement learning (RL) approaches are also highly effective for developing persuasive models, with such methods enabling more nuanced control over multiple persuasive dimensions through carefully designed reward functions.
Several studies have demonstrated the effectiveness of RL in enhancing specific persuasive qualities. Samad et al. (2022) augmented the PersuasionForGood dataset with emotion annotations and trained an agent using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm for more empathetically engaging and persuasive responses. Their reward function combined penalties for repetition and topical drift with incentives for emotional engagement and appropriate persuasive strategy use. Similarly, Mishra et al. (2022) refined emotional expression and politeness in model persuasive generations. They employed PPO with a reward that incorporates signals from persuasion, emotion, politeness strategy consistency, dialogue-coherence, and non-repetitiveness.
Tiwari et al. (2022a) applied RL to task-oriented dialogue agents, showing that persona-aware rewards could improve persuasive effectiveness toward shared goals. Similarly, Shi et al. (2021) also trained a GPT-2-based persuader with DialGAIL, which uses RL to address the issues of repetition, inconsistency, and task relevance. Most recently, (Hong et al., 2025) introduced an approach leveraging hindsight regeneration for training persuasive dialogue systems. They first simulate human-agent conversations and then retrospectively optimize agent responses based on the success of conversation outcomes, which is called hindsight regeneration. Finally, they employ offline RL on this data, resulting in improved persuasive performance.
Although reinforcement learning provides a powerful framework for shaping persuasive behavior, it remains underexplored. Future research should prioritize the development of more expressive and reliable reward functions that can capture subtle aspects of persuasive quality, such as ethical alignment, user receptivity, and long-term impact. In addition, there is substantial room to apply RL-based persuasion in many other domains. Methods like RLHF and RLAIF offer new opportunities for aligning persuasive generation with human values and feedback.

### v2

@@ 3.3. LLM Persuasiveness
Unlike the evaluation of argument persuasiveness, where arguments are predefined by humans, LLM persuasiveness evaluation focuses on model-generated content in both single-turn and multi-turn dialogues. In this survey, we refer to this setting as AI as Persuader, where the language model acts as the source of persuasion, aiming to influence a human’s beliefs, attitudes, or behaviors through generated arguments or dialogue. Leading companies such as OpenAI (OpenAI, 2024b), Anthropic (Durmus et al., 2024), and DeepMind (Phuong et al., 2024) have investigated the persuasiveness of their models to better understand the risks associated with deploying these highly conversational systems. However, evaluating the persuasive abilities of LLMs is not straightforward, as research explores a variety of evaluation methodologies and experimental setups.
@@ 3.3.1. Human Evaluation
Understanding the dynamics of AI to Human persuasion is crucial, as most proprietary models are designed for individual users. Given this, using human subjects to assess the persuasiveness of LLMs is a natural approach. Durmus et al. (2024) assessed model persuasiveness by measuring human agreement with a claim before and after exposure to persuasive arguments generated by Claude models. This single-turn setup reveals that larger models are generally more persuasive, with deceptive techniques emerging as particularly effective. In contrast, Phuong et al. (2024) introduced a suite of multi-turn evaluation benchmarks, including Money Talks, Charm Offensive, Hidden Agenda, and Web of Lies. These involve interactive persuasion tasks such as persuading users to donate, impersonating a friend, manipulating users to take suspicious actions on the computer, and promoting false beliefs. Unlike single-turn approaches, these multi-turn evaluations are expected to capture more dynamic and complex forms of persuasion.
@@
Similarly, Salvi et al. (2024) evaluated model persuasiveness through a debate game involving both human-human and human-LLM interactions. They found that participants were more likely to change their stance after engaging with GPT-4, regardless of whether the model had access to personal information. OpenAI (2024b) introduced two human evaluation benchmarks for assessing LLM persuasiveness. In the first setup, they used successful persuasion samples from the r/ChangeMyView subreddit as prompts, then asked LLMs to generate persuasive arguments in response. Human annotators scored both the original human arguments and the LLM-generated ones. The primary evaluation metric, referred to as the AI persuasiveness percentile relative to humans, measures the probability that a randomly chosen LLM-generated argument is rated as more persuasive than its human counterpart. The second benchmark, called Persuasion Parallel Generation, presents annotators with two arguments generated by different models for the same prompt. Annotators select the more persuasive argument, and the resulting win rates are used to compare model performance. Their results indicate that LLMs do not significantly outperform humans, and that newer models such as o1 do not show substantial gains in persuasiveness over earlier generations.
@@
Interestingly, similar lines of research have emerged within the social sciences. Huang and Wang (2023) conducted a meta-analysis examining whether artificial intelligence is more persuasive than humans. Using statistical methods, they compared persuasion outcomes—such as changes in perception, attitudes, intentions, and behaviors—between human-AI and human-human interactions. Their findings suggested that LLMs are about as persuasive as humans. Focusing more specifically on political communication, Goldstein et al. (2024) found that AI-generated propaganda is already as persuasive as real-world examples. Similarly, Bai et al. (2025) showed that AI performs on par with humans in crafting messages aimed at shaping public attitudes toward policy issues.
@@ 3.3.2. Automatic Evaluation
Singh et al. (2024) proposed PersuasionBench and PersuasionArena to evaluate LLMs’ persuasive abilities using tweet pairs from enterprise accounts that differ in engagement. They introduced simulative tasks (e.g., predicting or comparing tweet persuasiveness) and generative transsuasion tasks, defined as transforming non-persuasive content into persuasive content under controlled conditions to improve engagement. Their evaluations spanned standard NLP metrics, human judgments, and an Oracle language model. Another work in this direction, Pauli et al. (2025), trained a regression model to score the relative persuasiveness of LLM-generated rewrites, using human-annotated text pairs collected across diverse domains.
@@
Building on the conversational persuasion evaluation paradigm, OpenAI (2024b) proposed domain-specific scenarios to assess persuasive abilities. In MakeMeSay, the persuader is given a secret codeword and must steer the persuadee into saying it; in MakeMePay, the persuader adopts the role of a con artist attempting to elicit a payment. Similarly, Zhu et al. (2025) introduced a multi-agent bargaining task, where groups of LLM sellers attempt to persuade LLM buyers to purchase a product. Idziejczak et al. (2025) explored persuasion in a social deception game inspired by Among Us, where LLMs played as either crewmates or impostors. Persuasive behavior emerged through dialogue, which was later annotated by another LLM using predefined persuasion strategies to assess how models influenced one another within the game.
While early findings suggest that LLMs can be as persuasive as humans—and in some cases more so, particularly as they improve in reasoning, personalization, and strategic communication, the underlying mechanisms driving this persuasiveness remain poorly understood. Current research has revealed that different evaluation setups may yield inconsistent or even contradictory results, with some models outperforming others in one framework while falling short in another. LLM-based evaluations may also capture persuasion between models rather than effects on humans, risking circularity. This highlights the complex, multidimensional nature of persuasive ability and raises important concerns about the increasing influence of ”smarter” models on beliefs and behaviors. As a result, developing a unified and trustworthy evaluation framework that integrates diverse test cases and assesses persuasive skills across multiple dimensions is a pressing research challenge. Such a framework would not only support more self-consistent and interpretable evaluations but also help explain why certain models excel in specific contexts while lacking in others, ultimately informing safer and more responsible deployment of persuasive AI systems.
@@ 4. Generating Persuasion
AI-driven systems designed to generate persuasive content are becoming increasingly powerful and widespread. Persuasion can play a pivotal role in various domains, including advertisements, healthcare promotion, recommendation systems, political campaigns, and more. As organizations and individuals seek to increase their influence, the need for generating persuasive content has grown significantly, especially with the advent of LLMs. This section explores prior work on enhancing persuasive capabilities, examines key influencing factors, and highlights applications of persuasion in settings like negotiation and debate.
@@ 4.1. Methods
Section 2.2.1 introduced taxonomies of persuasive strategies that LLMs can adopt to enhance the effectiveness of their outputs. Building on this foundation, recent work has focused on developing methods to improve LLMs’ ability to generate persuasive content. In this subsection, we review key approaches proposed to enhance the persuasive capabilities of LLMs, including prompting techniques, incorporation of external knowledge, fine-tuning, and reinforcement learning (See Table 3). Beyond neural approaches, structured reasoning methods from computational argumentation offer an alternative way to model persuasive generation. Argumentation-based chatbots have proven particularly effective in persuasion, where explicit reasoning and dialogue structure support goal-directed interaction (Castagna et al., 2024). Integrating such methods with LLMs could enable hybrid systems that combine the fluency of neural models with the interpretability and deliberative rigor of formal argumentation.
@@
(Tiwari et al., 2022a) Personalization PPMD (Tiwari et al., 2022a) Selects a persuasive strategy based on the user profile
@@
(Tiwari et al., 2022b) Personalization DevPVA (Tiwari et al., 2022b) Rewards persuasion based on persona of the user
@@ 4.1.1. Prompting
Another approach is to prompt LLMs with specific persuasive strategies. Durmus et al. (2024) examined how different types of prompts affect the persuasiveness of model outputs by experimenting with four prompting strategies: (1) prompting the model to construct compelling arguments, (2) adopting the role of an expert and using rhetorical techniques such as pathos, logos, and ethos, (3) encouraging logical reasoning, and (4) permitting deception, including fabricating evidence, to boost persuasive impact. Their findings showed that while effectiveness varies across model architectures, prompts that encourage logical reasoning or allow for deception tend to produce more persuasive arguments overall. Along similar lines, Zeng et al. (2024) designed a comprehensive taxonomy of persuasion strategies, then used these strategies to generate persuasive adversarial prompts which proved to be remarkably effective in jailbreaking LLMs. Using persuasive strategies has also demonstrated effectiveness in misinformation propagation, where persuasive strategies effectively manipulate LLMs into endorsing false information (Xu et al., 2024a).
@@ 4.1.2. Incorporating External Information
While prompting offers a good baseline for persuasive generation, incorporating external information can substantially enhance persuasiveness by introducing additional dimensions such as personalization and factual grounding. Models can produce more tailored and credible persuasive content by leveraging contextual knowledge, including user profiles, background information, or external evidence. Effective persuasion involves not only implementing persuasive strategies but also selecting techniques that align with the characteristics of the recipient. Lukin et al. (2017) and Wang et al. (2019) showed that accounting for personality traits is critical, as individuals with different personality types tend to respond differently to specific persuasive strategies. Building on this insight, Jin et al. (2024) proposed a method where LLMs begin by identifying and summarizing the user’s intent before reasoning about the most suitable strategy to apply. This process enables a more adaptive and responsive persuasive system.
Incorporating personal information about the user extends beyond just strategy selection. Personalization based on users’ psychological traits, such as personality, political ideology, moral foundations, and behavioral patterns, has been shown to significantly enhance persuasive impact (Matz et al., 2024; Kaptein et al., 2015). This approach is effective across various domains, including targeted advertisements and political messaging around climate action (Matz et al., 2024; Simchon et al., 2024; Meguellati et al., 2024). Salvi et al. (2024) showed that providing LLMs with personal information during human interaction increases their persuasiveness. Research by Ruiz-Dolz et al. (2024) demonstrated that user-specific persuasive policies significantly improve the persuasiveness of argument-based systems. Zhang and Zhou (2025) emphasized the importance of considering the user’s mental state when constructing persuasive arguments. Tiwari et al. (Tiwari et al., 2022a, 2023) showed that combining dialogue context with user profile information provides further persuasive benefits. In parallel, other research highlighted the value of counterarguments that directly engage with the user’s concerns or beliefs (Cima et al., 2024; Hunter et al., 2019).
@@ 4.1.3. Finetuning
Finetuning LLMs on persuasive datasets offers a powerful way to enhance their persuasive capabilities beyond what prompting alone can achieve. This approach allows models to internalize persuasive patterns and strategies at a more fundamental level. In the negotiation domain, Lewis et al. (2017) trained a model that is based on four recurrent neural networks on negotiation datasets, and demonstrated significant improvements in negotiation ability. Extending finetuning to emotional support contexts, Liu et al. (2021) introduced the Emotion Support Conversation dataset (ESConv) and showed that finetuning BlenderBot (Roller et al., 2021) and DialoGPT (Zhang et al., 2020) on this data enhances models’ ability to provide empathetic responses and sustain engagement.
Chen et al. (2022) finetuned BART-large (Lewis et al., 2020) on the PersuasionForGood dataset (Wang et al., 2019), achieving significant gains in persuasive quality. More recently, Jin et al. (2024) developed PersuGPT by training LLaMa-2-13B-Chat on the DailyPersuasion dataset, which includes rich reasoning traces. Combining conventional finetuning with Direct Preference Optimization (DPO) (Rafailov et al., 2023), their model learns to infer user intent, select appropriate persuasive strategies, and anticipate future user responses and rewards. Singh et al. (2024) proposed an instruction-based finetuning approach to enhance LLM persuasiveness. They finetuned Vicuna-1.5 13B to rewrite inputs to be more persuasive and showed that the resulting model outperforms the base version and can rival much larger LLMs in persuasiveness. Finetuning presents strong potential for developing more persuasive language models, yet remains relatively underexplored. With more carefully curated training data and domain-specific finetuning, future work can yield models that are better aligned with persuasive goals across a range of applications, including healthcare, education, and advertising.
@@ 4.1.4. Reinforcement Learning
Reinforcement learning (RL) approaches are also highly effective for developing persuasive models, enabling more nuanced control over multiple persuasive dimensions through carefully designed reward functions. Several studies have shown the effectiveness of RL in enhancing persuasive qualities. Samad et al. (2022) augmented the PersuasionForGood dataset with emotion annotations and trained an agent using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm for more empathetically engaging and persuasive responses. Their reward function combined penalties for repetition and topical drift with incentives for emotional engagement and appropriate persuasive strategy use. Similarly, Mishra et al. (2022) refined emotional expression and politeness in model persuasive generations. They employed PPO with a reward that incorporates signals from persuasion, emotion, politeness strategy consistency, dialogue-coherence, and non-repetitiveness.
Tiwari et al. (2022b) applied RL to task-oriented dialogue agents, showing that persona-aware rewards could improve persuasive effectiveness toward shared goals. Similarly, Shi et al. (2021) also trained a GPT-2-based persuader with DialGAIL, which uses RL to address the issues of repetition, inconsistency, and task relevance. Most recently, (Hong et al., 2025) introduced an approach leveraging hindsight regeneration for training persuasive dialogue systems. They first simulate human-agent conversations and then retrospectively optimize agent responses based on the success of conversation outcomes, which is called hindsight regeneration. Finally, they employ offline RL on this data, resulting in improved persuasive performance. Although RL provides a powerful framework for shaping persuasive behavior, it remains underexplored. Future research should prioritize the development of more expressive and reliable reward functions that can capture subtle aspects of persuasive quality, such as ethical alignment, user receptivity, and long-term impact. In addition, there is substantial room to apply RL-based persuasion in many other domains. Methods like RLHF and RLAIF offer new opportunities for aligning persuasive generation with human values and feedback.